# FT: Unlocking the Power of Code Instruction Tuning by Simply Merging Upcycled Mixture-of-Experts

### v1

@@ 4.1 Experimental Setup
Training. We use DeepSeek-Coder-Base 1.3B Guo et al. (2024) as the main base code LLM. evol-codealpaca-v1, an open-source Evol-Instruct Luo et al. (2023) dataset containing 110K samples, is used as our instruction dataset. MoE, our MoE model upcycled from the base model, is implemented following Llama-MoE LLaMA-MoE Team (2023). It is constructed with 8 experts in one expert layer and the top 6 experts111 6 is the best-performing number of activated experts per our HumanEval+ experiments using top experts. are activated for each token, including one shared expert. As such, we denote the model size of MoE as 81.3B. Other hyperparameter settings are detailed in Appendix A.1. We finally obtain FT by using the learned mixing coefficients to merge MoE layers inside MoE as normal FFN layers. Note that FT is the final instruction-tuned LLM we produce, while MoE is only an intermediate product of FT framework.
Baselines. To study the effectiveness of FT, we build a baseline model, namely SFT, by directly performing SFT for DeepSeek-Coder-Base 1.3B on evol-codealpaca-v1. To compare FT with EWA Huang et al. (2023), we also implement a baseline EWA and instruction-tune it using the same hyperparameter setting as SFT, which is described in Appendix A.1. More implementation details of EWA can be seen in Appendix A.2. Furthermore, we incorporate multiple small open-source models (<3B) as our baselines, including DeepSeek-Coder-Base 1.3B, DeepSeek-Coder-Instruct 1.3B Guo et al. (2024), Phi-2 2.7B, and stable-code 3B Pinnaparaju et al. (2024).
@@ 4.2 Python Text-to-Code Generation
HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks are the two most widely-used collections of Python code generation tasks. We further employ HumanEval+ and MBPP+, which use more tests automatically generated by EvalPlus Liu et al. (2023) for more rigorous evaluation. We leave the details in Appendix A.3.
Table 1 shows the pass@1 results of different LLMs. FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which makes it the new state-of-the-art small code LLM (<3B). We can also observe that FT has a clear improvement over the SFT on both benchmarks, with 13% and 2% improvement on HumanEval+ and MBPP+ respectively, while EWA even performs worse than SFT on MBPP(+). FT also outperforms EWA on both benchmarks. Surprisingly, FT even surpasses MoE on HumanEval and HumanEval+, despite only using around parameters and around computations, which showcases the effectiveness of our simple learnable merging technique. Appendix A.4 further demonstrates the statistical significance of the improvements brought by FT.
@@ 4.3 Multilingual Code Generation
We use MultiPL-E Cassano et al. (2022), a multi-programming benchmark that supports 18 programming languages in addition to Python, to evaluate the multilingual ability and generalizability of FT. Among these, we choose 6 representative programming for their distinct language features: Java, JavaScript, C++, PHP, Swift, and Rust, following Wei et al. (2023). Table 2 shows, among all 1.3B models, FT achieves the best average multilingual performance and performs the best on 5 (out of 6) individual programming languages, overall largely improving SFT which uses standard SFT. Notably, the overall performance of EWA is on par with SFT, indicating that EWA may not improve SFT on multilingual coding. Appendix A.5 further studies whether each expert in MoE specializes differently in these programming languages.
@@ 4.4 Code Generation for Data Science
The DS-1000 dataset Lai et al. (2022) is a collection of 1000 realistic data science coding problems ranging from 7 popular data science libraries in Python, including Matplotlib (plt), NumPy (np), Pandas (pd), SciPy (scp), Scikit-Learn (sk), PyTorch (py), and TensorFlow (tf). We evaluate FT on DS-1000 to understand its effectiveness for practical data science engineering. We follow the evaluation setting of prior works Guo et al. (2024); Wei et al. (2023). In Table 3, FT achieves the best overall performance among all the evaluated 1.3B models. Specifically, FT consistently surpasses SFT among all the seven studied libraries and also outperforms EWA in general.
@@ 5.1 Effect of Shared Expert with Routing Weight Normalization
We demonstrate the importance of the shared expert of FT by comparing its performance with the sparse upcycling Komatsuzaki et al. (2023) baseline that does not employ any shared expert. As shown in Table 4, the performance of the original sparse upcycling (with the "- Shared Expert" label) drops greatly compared with MoE. Notably, the sparse upcycling model even performs worse than SFT on HumanEval+, showing its ineffectiveness for instruction tuning.
While the shared expert setting is also employed in most recent works Dai et al. (2024); Gou et al. (2024), their routing strategy will cause performance degradation due to the scale mismatch between the outputs of the upcycled MoE layer and the original FFN layer. To understand the importance of routing weight normalization, we conduct an ablation by excluding it from FT. Table 4 shows that, after removing routing weight normalization, the performance substantially decreases, despite being still better than the original sparse upcycling that does not use the shared expert setting.
@@ 5.2 Effect of Merging Strategy
In this section, we demonstrate the effectiveness of our learnable merging technique by comparing it with (1) directly merging experts with initialized mixing coefficients, and (2) the learnable merging technique without the shared rate setting, which is the same setting as the learned soup in Model Soups Wortsman et al. (2022) and is described in Eq. (3) and Eq. (4). Specifically, we initialize the learnable mixing coefficient of the shared expert as 0.75 and that of the other 7 normal experts as for fair comparison. As shown in Table 5, trained mixing coefficients outperform the initialized mixing coefficients for merging. Furthermore, removing the shared rate setting will largely degrade the performance of FT on both HumanEval and HumanEval+, demonstrating its importance.
| Model | HumanEval | HumenEval+ | |
| SFT | - | 61.6 | 57.3 |
| FT | 0.00 | 62.8 | 59.8 |
| 0.25 | 64.6 | 61.0 | |
| 0.50 | 65.9 | 62.8 | |
| 0.75 | 67.1 | 64.6 | |
| 1.00 | 63.4 | 60.4 |
The performance of the final merged dense model drops significantly when the shared expert rate grows from 0.75 to 1.00, showing that specific knowledge learned by other experts is also irreplaceable and ignoring them will lead to a significant performance drop.
5.3 Effect of Base Code LLM
In this section, we demonstrate that the effectiveness of FT is not dependent on the choice of base code LLMs. To show this, we conduct an ablation experiment by applying FT to stable-code 3B Pinnaparaju et al. (2024), whose architecture is different from DeepSeek-Coder-Base 1.3B Guo et al. (2024), and see whether it can still improve its performance. Hyperparameter settings are detailed in Appendix A.6. As is shown in Table 7, FT significantly improves SFT by 10% on HumanEval and 11% on HumanEval+ respectively. Furthermore, FT consistently boosts the performance of MoE while only using parameters and computations. These results show that the effectiveness of FT does not depend on any specific choice of base code LLMs, demonstrating the generalizability of FT across different model architectures.
@@ 6 Discussion
6.1 Training Overhead Analysis
Compared with SFT, FT will inevitably introduce additional overhead in the training process because FT needs to fine-tune the upcycled MoE model while the normal SFT technique only needs to fine-tune the original dense model. To better understand the effect of such overhead, we conduct an experiment where we use the same training budget (i.e., the same GPU hours) instead of the same training steps for the normal SFT baseline. As shown in Table 8, although sharing the same training budget as FT, the performance of SFT is still significantly worse than that of FT, demonstrating the ability of FT to unlock the power of code instruction tuning using the same training budget.
| Model | HumanEval | HumanEval+ | ||
61.6 | 57.3 | ||
62.2 | 57.3 | ||
| FT | 67.1 | 64.6 |
6.2 Generalizability for General Tasks
@@
In this section, we demonstrate that FT can improve the performance of LLMs on general tasks across different domains by applying FT to general instruction tuning. We use TinyLlama 1.1B Zhang et al. (2024) as the base model and use evol-instruct-70k Xu et al. (2023) as the training dataset for general instruction tuning. Following existing work Zhang et al. (2024), we use MMLU Hendrycks et al. (2021) with the 5-shot setting as our evaluation benchmark to showcase the general performance of LLMs. Hyperparameter settings are detailed in Appendix A.7. As shown in Table 9, overall, FT improves SFT by 5% on MMLU, demonstrating the generalizable effectiveness of FT for general instruction tuning.
@@ 6.3 Preliminary Theoretical Explanation
Step 1: Upcycling. According to the scaling law Kaplan et al. (2020), the upcycled MoE model performs better than the normal SFT dense model due to more trainable parameters.
Step 2: Merging. We consider a simplified variant of FT, where the upcycled MoE model (e.g., MoE) can be viewed as the ensembling of two dense models and the merged dense model (e.g., FT) can be viewed as the merging of the same two dense models; see Appendix A.8 for more details. As such, we can directly apply the theoretical analyzing process in Section 4 of Wortsman et al. (2022) to analyze the performance difference between the upcycled MoE model and the merged dense model, which is initially designed to analyze the performance difference between model ensembling and model merging. According to Wortsman et al. (2022), the convexity of the loss can help the merged dense model achieve a similar expected loss as that of the upcycled MoE model.
Overall, the Upcycling step improves the performance with more trainable parameters, while the Merging step maintains the upcycled MoE-level performance with only dense-model compute. Consequently, we provide a preliminary theoretical explanation for the effectiveness of FT.
@@ 7 Conclusion
This paper introduces FT to unlock the power of code instruction tuning by simply merging upcycled MoE. Similar to SFT, FT starts with a dense LLM and produces a fine-tuned dense LLM with the exact size and model structure. Yet, FT improves SFT by upcycling the pre-trained dense LLM to an MoE model for fine-tuning, after which we compile the MoE model back to an efficient dense LLM with a learnable merging mechanism. As such, we unleash the performance limit of instruction tuning without any additional inference overhead. Using the same dataset, FT improves SFT on a variety of benchmarks, including HumanEval(+), MBPP(+), MultiPL-E, and DS-1000, from 2% to 13%. By applying FT to DeepSeek-Coder-Base 1.3B, we create the next state-of-the-art small (<3B) LLM for code. The ultimate dense LLM produced by FT preserves or even outperforms the full upcycled MoE which uses parameters as much as our final dense LLM. FT is fully orthogonal to the existing instruction tuners such as Evol-Instruct and OSS-Instruct, opening a new dimension to maximal code instruction tuning.
@@
While FT has proven to be effective through extensive experiments in the paper, we apply our technique to LLMs with no more than 3B parameters due to resource constraints. This limitation hinders our ability to showcase the impact of FT on larger models. In addition, to balance the general knowledge in the shared expert and the specific knowledge in other normal experts, we introduce a hyperparameter in the merging process of FT, which might slightly increase the efforts for hyperparameter search. It would be interesting to explore other hyperparameter-free techniques to tackle this challenge in the future. Furthermore, while FT has been empirically proven powerful, it would be interesting to provide a theoretical explanation for its strong performance.
@@
A.1 Hyperparameter Settings
We use a batch size of 64 and a learning rare of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 500 warmup steps, following the implementation of previous work Wei et al. (2023). We further use a batch size of 64, a shared expert rate of 0.75, and a learning rare of 1e-5 with a linear schedule to fine-tune the learnable mixing coefficients for each of the experts in the instruction-tuned MoE on the instruction dataset for 1 epoch with 125 warmup steps. Detailedly, we use Softmax to keep the sum of the mixing coefficients of the other 7 normal experts as 0.25. For SFT and EWA, we use the same hyperparameter setting as FT, where the batch size is 64 and the learning rate is 5e-5 with a linear scheduler. Because FT is trained for 4 epochs during upcycling and 1 epoch during merging, for a fair comparison, we train SFT and EWA for 5 (= 4 + 1) epochs with 625 warmup steps.
@@ A.2 Implementation details of EWA
Because EWA Huang et al. (2023) does not release their implementation, we implemented EWA by ourselves, including constant schedule and linear schedule. We use a share rate of 0.3, following the original setting of EWA. While EWA with the constant schedule achieves reasonable performance in our evaluation, the training loss of EWA with the linear schedule becomes very unstable, as is shown in Figure 3, and thus cannot achieve reasonable performance. As a result, we report the results of EWA with the constant schedule in Section 4.
A.3 Details of HumanEval and MBPP
@@ A.4 Statistical Significance Analysis
In this section, we show that improvements brought by FT are statistically significant. In our main experiments, we follow prior works Wei et al. (2023); Lozhkov et al. (2024) to conduct experiments on HumanEval (+) using greedy decoding. To demonstrate the statistical significance of our improvements, we change our setting from greedy decoding to sampling. In detail, to conduct one experiment on HumanEval (+), the model will sample one solution for each problem in HumanEval (+) with top = 0.95 and temperature = 0.8, which is the same setting used in prior works Liu et al. (2023); Chen et al. (2021).
@@
Following prior work Liu et al. (2023), we repeat this experiment 200 times for three techniques: FT, EWA, and SFT. EWA is included because it is the best-performing baseline in our main experiment. We first compute their average pass@1 performance in these 200 experiments. As is shown in Table 10, FT outperforms both EWA and SFT clearly.
@@
Furthermore, we use the Wilcoxon signed-rank test Wilcoxon (1945); Dror et al. (2018), a widely used statistical test, to check if the improvements brought by FT are statistically significant. As shown in Table 11, the -values for both FT vs. EWA and FT vs. SFT are much smaller than both 0.0025 (the significance level recommended for NLP work by Søgaard et al. (2014)) and 0.05 (the most common significance level), demonstrating the statistical significance of the improvements brought by FT.
A.5 Analysis on Expert Specialization

### v2

@@ 4.1 Experimental Setup
Training. DeepSeek-Coder-Base 1.3B Guo et al. (2024) is used as our main base code LLM. evol-codealpaca-v1, an open-source Evol-Instruct Luo et al. (2023) dataset containing 110K samples, is used as our code instruction dataset. MoE, our MoE model upcycled from the base model, is implemented following Llama-MoE LLaMA-MoE Team (2023). It is constructed with 8 experts in one MoE layer and the top 6 experts111 6 is the best-performing number of activated experts per our HumanEval+ experiments using top experts. are activated for each token, including one shared expert. As such, we denote the model size of MoE as 81.3B. Other training settings are detailed in Appendix A.1. We further obtain FT by using the learned mixing coefficients to compile MoE layers inside MoE to normal FFN layers. Note that FT is the final instruction-tuned code LLM we output, while MoE is only an intermediate product of our FT framework.
Baselines. To study the effectiveness of FT, we build a baseline model, namely SFT, by directly performing SFT for DeepSeek-Coder-Base 1.3B on evol-codealpaca-v1. To compare FT with EWA Huang et al. (2023), we also implement a baseline EWA and instruction-tune it using the same hyperparameter setting as SFT, which is described in Appendix A.1. More implementation details of EWA are described in Appendix A.2. Furthermore, we incorporate multiple tiny open-source LLMs (<3B) as our baselines, including DeepSeek-Coder-Base 1.3B, DeepSeek-Coder-Instruct 1.3B Guo et al. (2024), Phi-2 2.7B, and stable-code 3B Pinnaparaju et al. (2024).
@@ 4.2 Python Text-to-Code Generation
HumanEval Chen et al. (2021) and MBPP Austin et al. (2021) benchmarks are the two most widely-used collections of Python code generation tasks. We further employ HumanEval+ and MBPP+, which use more tests automatically generated by EvalPlus Liu et al. (2023) for more rigorous evaluation. We leave the detailed description of HumanEval(+) and MBPP(+) in Appendix A.3.
Table 1 shows the pass@1 results of different LLMs. FT achieves 67.1 pass@1 on HumanEval and 64.6 pass@1 on HumanEval+, which makes it the new state-of-the-art tiny code LLM (<3B). We can also observe that FT has a clear improvement over the SFT on both benchmarks, with 13% and 2% improvement on HumanEval+ and MBPP+ respectively. In contrast, EWA not only underperforms FT on both benchmarks, but also fails to improve SFT on MBPP(+). Surprisingly, FT even surpasses MoE on HumanEval and HumanEval+, despite only using around parameters and around computations, which showcases the effectiveness of our simple learnable merging technique. More comprehensive experiments in Appendix A.4 demonstrate the statistical significance of the improvements brought by FT. Furthermore, while FT will inevitably introduce training overhead, our experiment in Appendix A.5 shows that FT still significantly outperforms SFT using the same training budget, demonstrating the ability of FT to unlock the power of code instruction tuning.
@@ 4.3 Multilingual Code Generation
We use MultiPL-E Cassano et al. (2022), a multi-programming benchmark that supports 18 programming languages in addition to Python, to evaluate the multilingual ability and generalizability of FT. Following previous work Wei et al. (2023), we choose six representative programming languages in our evaluation for their distinct language features: Java, JavaScript, C++, PHP, Swift, and Rust. Table 2 shows that, among all 1.3B models, FT achieves the best average multilingual performance and performs the best on five (out of six) programming languages and overall largely improves SFT which uses standard SFT. Notably, the overall performance of EWA is on par with SFT, indicating that EWA fails to improve SFT on multilingual coding. Appendix A.6 further studies whether each expert in MoE specializes differently in different programming languages.
@@ 4.4 Code Generation for Data Science
The DS-1000 dataset Lai et al. (2022) is a collection of 1000 realistic data science coding problems ranging from seven popular data science libraries in Python, including Matplotlib (plt), NumPy (np), Pandas (pd), SciPy (scp), Scikit-Learn (sk), PyTorch (py), and TensorFlow (tf). We evaluate FT on DS-1000 to understand its effectiveness for practical data science engineering. We follow the evaluation setting of prior works Guo et al. (2024); Wei et al. (2023). Table 3 shows that FT achieves the best overall performance among all the evaluated 1.3B models. Specifically, FT consistently surpasses SFT among all the seven studied libraries and outperforms EWA in general.
@@ 5.1 Effect of Shared Expert with Routing Weight Normalization
We demonstrate the importance of the shared expert of FT by comparing its performance with the original sparse upcycling Komatsuzaki et al. (2023) baseline that does not employ any shared expert. As shown in Table 4, the performance of the original sparse upcycling (with the "- Shared Expert" label) drops greatly compared with MoE. Notably, the sparse upcycling model performs even worse than SFT on HumanEval+, indicating its ineffectiveness for instruction tuning.
While the shared expert setting is also employed in most recent works Dai et al. (2024); Gou et al. (2024), their routing strategy will cause performance degradation due to the scale mismatch problem, which is handled by the routing weight normalization design in FT. To demonstrate its importance, we conduct an ablation experiment by excluding it from FT. Table 4 shows that, after removing routing weight normalization, the performance substantially decreases, despite still performing better than the original sparse upcycling that does not use the shared expert setting.
@@ 5.2 Effect of Merging Strategy
In this section, we demonstrate the effectiveness of our learnable merging technique by comparing it with (1) directly merging experts with initialized mixing coefficients, and (2) the learnable merging technique without the shared expert rate setting, which is the same setting as the learned soup in Model Soups Wortsman et al. (2022) and is described in Eq. (3) and Eq. (4). Specifically, we initialize the learnable mixing coefficient of the shared expert as 0.75 and that of the other 7 normal experts as for a fair comparison. As shown in Table 5, trained mixing coefficients outperform the initialized mixing coefficients for merging. Furthermore, removing the shared rate setting will degrade the performance of FT on both HumanEval and HumanEval+, demonstrating its importance. An ablation study on the shared expert rate in Appendix A.7 further shows that (1) FT consistently outperforms SFT regardless of their shared expert rate, and (2) both the general knowledge learned in the shared expert and the specific knowledge learned in other experts are important and integral for better performance after merging.
5.3 Effect of Code LLM Choice
To show that the effectiveness of FT is not dependent on any specific code LLMs, we apply FT to stable-code 3B Pinnaparaju et al. (2024), whose architecture is different from DeepSeek-Coder-Base 1.3B Guo et al. (2024), to study whether FT can still improve the performance of this new model. The training settings are detailed in Appendix A.8. As shown in Table 6, FT significantly improves SFT by 10% on HumanEval and 11% on HumanEval+ respectively. Furthermore, FT consistently boosts the performance of MoE while only using parameters and inference computations. These results show that the effectiveness of FT is generalizable across different code LLMs.
@@ 6 Discussion
| Model | HumanEval | HumanEval+ |
| SFT | 77.4 | 70.7 |
| MoE | 81.1 | 75.6 |
| FT | 81.7 | 76.8 |
6.1 Scaling up FT to 7B Scale
@@
We scale up FT to 7B-level code LLMs by applying it to DeepSeek-Coder-Base 6.7B Guo et al. (2024). The training settings are detailed in Appendix A.9. As shown in Table 7, FT significantly improves SFT by 6% on HumanEval and 9% on HumanEval+ respectively. Moreover, FT further boosts the performance of MoE with only parameters and computations during inference! These promising results demonstrate the consistent effectiveness of FT on 7B-level code LLMs.
6.2 Generalizability for General Tasks
To demonstrate that FT can also improve the performance of LLMs on general tasks across different domains, we apply FT to general instruction tuning. We use TinyLlama 1.1B Zhang et al. (2024) as the base model and use evol-instruct-70k Xu et al. (2023) as the training dataset for general instruction tuning. Following existing work Zhang et al. (2024), we use MMLU Hendrycks et al. (2021) with the 5-shot setting as our evaluation benchmark to evaluate the general performance of instruction-tuned LLMs. More training settings are detailed in Appendix A.10. As shown in Table 8, FT improves SFT by 5% on MMLU in general, demonstrating the generalizable effectiveness of FT for general instruction tuning.
@@ 6.3 Preliminary Theoretical Explanation
Step 1: Upcycling. According to the scaling laws Kaplan et al. (2020), the upcycled MoE model performs better than the normal SFT dense model due to more trainable parameters.
Step 2: Merging. We consider a simplified variant of FT, where the upcycled MoE model (e.g., MoE) can be viewed as the ensembling of two dense models and the merged dense model (e.g., FT) can be viewed as the merging of the same two dense models. More details are included in Appendix A.11. As such, we can directly apply the theoretical analyzing process in Section 4 of Model Soups Wortsman et al. (2022) to analyze the performance difference between the upcycled MoE model and the merged dense model, which is initially designed to analyze the performance difference between model ensembling and model merging. According to the analysis Wortsman et al. (2022), the convexity of the loss can help the merged dense model achieve a similar expected loss as that of the upcycled MoE model.
Overall, our preliminary theoretical explanation shows that (1) the Upcycling step improves the performance with more trainable parameters, and (2) the Merging step can provably maintain the performance of the aforementioned simplified MoE model with only dense-model compute.
@@ 7 Conclusion
This paper introduces FT to unlock the power of code instruction tuning by simply merging upcycled MoE. Similar to SFT, FT starts with a dense LLM and outputs a fine-tuned dense LLM with the exact size and model structure. Yet, FT improves SFT by upcycling the pre-trained dense LLM to an MoE model for fine-tuning, after which we compile the MoE model back to an efficient dense LLM with a learnable merging mechanism. As such, we unleash the performance limit of instruction tuning without any additional inference overhead. Using the same training dataset, FT outperforms SFT on a variety of benchmarks, including HumanEval(+), MBPP(+), MultiPL-E, and DS-1000, from 2% to 13%. By applying FT to DeepSeek-Coder-Base 1.3B, we create the new state-of-the-art tiny code LLM (<3B). The final dense LLM produced by FT preserves or even outperforms the full upcycled MoE which uses parameters as much as our final dense LLM. FT is fully orthogonal to the existing instruction tuners such as Evol-Instruct and OSS-Instruct, opening a new dimension for code instruction tuning.
@@
To balance the general knowledge in the shared expert and the specific knowledge in other normal experts, we introduce a hyperparameter in the merging process of FT, which might slightly increase the efforts for hyperparameter search. It would be interesting to explore other hyperparameter-free techniques to tackle this challenge. Furthermore, while we have provided a preliminary theoretical explanation for the strong empirical performance of FT, it would be interesting to provide a complete theoretical explanation in the future.
Acknowledgement
We extend our special thanks to Terry Yue Zhuo for his assistance with the scale-up experiments on DeepSeek-Coder-Base 6.7B (Section 6.1) after our submission. His contributions are good enough to merit authorship; however, due to the policy of ACL 2024, post-submission authorship changes are not permitted. As a result, we have included him in the author list of our arXiv version. We also thank Sea AI Lab and Dr. Qian Liu for their valuable feedback and computing resource assistance. We appreciate all the reviewers for their insightful comments. This work was partially supported by NSF grant CCF-2131943, as well as Kwai Inc.
@@ Acknowledgement
A.1 Training Settings for DeepSeek-Coder-Base 1.3B
We use a batch size of 64 and a learning rate of 5e-5 with a linear scheduler to fine-tune MoE for 4 epochs with 500 warmup steps, following the implementation of previous work Wei et al. (2023). We further use a batch size of 64, a shared expert rate of 0.75, and a learning rate of 1e-5 with a linear schedule to fine-tune the learnable mixing coefficients for experts in the instruction-tuned MoE on the same instruction-tuning dataset for 1 epoch with 125 warmup steps. Detailedly, we use Softmax to keep the sum of the mixing coefficients of the other 7 normal experts as 0.25. For SFT and EWA, we use the same hyperparameter setting as FT, where the batch size is 64 and the learning rate is 5e-5 with a linear scheduler. Because FT is trained for 4 epochs during upcycling and 1 epoch during merging, for a fair comparison, we train SFT and EWA for 5 (= 4 + 1) epochs with 625 warmup steps.
@@ A.2 Implementation details of EWA
Because EWA Huang et al. (2023) does not release their implementation, we implement EWA by ourselves, including the constant schedule setting and the linear schedule setting. We use a share rate of 0.3, following the original setting of EWA. While EWA with the constant schedule setting achieves reasonable performance in our evaluation, the training loss of EWA with the linear schedule setting is unstable during instruction tuning, as shown in Figure 3, and thus cannot achieve reasonable performance. As a result, we report the results of EWA with the constant schedule setting in Section 4.
A.3 Details of HumanEval(+) and MBPP(+)
@@ A.4 Statistical Significance Analysis
In our main experiments, we follow prior works Wei et al. (2023); Lozhkov et al. (2024) to conduct experiments on HumanEval(+) using greedy decoding. To demonstrate the statistical significance of our improvements, we change our setting from greedy decoding to sampling. In detail, to conduct one experiment on HumanEval(+), the model will sample one solution for each problem in HumanEval(+) with top = 0.95 and temperature = 0.8, which is the same setting used in prior works Liu et al. (2023); Chen et al. (2021).
@@
Following prior work Liu et al. (2023), we repeat this experiment 200 times for three techniques: FT, EWA, and SFT. EWA is included because it is the best-performing baseline in our main experiment. We first compute their average pass@1 performance in these 200 experiments. As is shown in Table 9, FT outperforms both EWA and SFT.
@@
Furthermore, we use the Wilcoxon signed-rank test Wilcoxon (1945); Dror et al. (2018), a widely used statistical test, to check if the improvements brought by FT are statistically significant. As shown in Table 10, the -values for both FT vs. EWA and FT vs. SFT are much smaller than both 0.0025 (the significance level recommended for NLP work Søgaard et al. (2014)) and 0.05 (the most common significance level), demonstrating the statistical significance of the improvements brought by FT.
A.5 Training Overhead Analysis
| Model | HumanEval | HumanEval+ | ||
61.6 | 57.3 | ||
62.2 | 57.3 | ||
| FT | 67.1 | 64.6 |
Compared with SFT, FT will inevitably introduce additional overhead in the training process because FT needs to fine-tune the upcycled MoE model, which contains more parameters than the original dense model and thus requires more computation. In contrast, the normal SFT technique only needs to fine-tune the original dense model. To better understand the effect of such overhead, we conduct an experiment using the same training budget (i.e., the same GPU hours) instead of the same training steps for the normal SFT baseline. As shown in Table 11, although sharing the same training budget as FT, the performance of SFT is still significantly worse than that of FT, demonstrating the ability of FT to unlock the power of code instruction tuning using the same training budget.
A.6 Expert Specialization Analysis